8(4u - 1)-12u = 11(2u - 6)
Answer by umail08(87)

8(4u - 1)-12u = 11(2u - 6)
32u – 8 – 12u = 22u – 66
20u – 8 = 22u – 66
20u – 22u = -66 + 8
-2u = -58
u = 29